April 13, 2022 Calorimeter

From GlueXWiki
Jump to: navigation, search

Meeting Day/Time: every other WEDNESDAY at 11:00 a.m. JLab time

Connection instructions

Use ZoomGov: https://jlab-org.zoomgov.com/j/1608552106?pwd=UE1nZFIzUUZHVjJEVGJXTGdUYTV3UT09.
Our meeting ID number is 160 855 2106. Click "Expand" for more details.

  • Meeting ID: 160 855 2106 Passcode: 944454
  • One tap mobile
       +16692545252,,1608552106# US (San Jose)
       +16468287666,,1608552106# US (New York)
  • Dial by your location
       +1 669 254 5252 US (San Jose)
       +1 646 828 7666 US (New York)
       +1 551 285 1373 US
       +1 669 216 1590 US (San Jose)
       833 568 8864 US Toll-free (US West) (US East)
       Meeting ID: 160 855 2106
       Passcode: 944454


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book
  3. CCAL HDCCAL log book
  4. Calibrations: 2020 Data Production; RunPeriod-2019-11 Validation; Offline Monitoring Data Validation
  5. Photon Reconstruction Task Force
  6. BCAL calibration links, 2020, PrimEx, SRC

Goals for Calorimetry Group

  1. Determine preliminary photon reconstruction efficiencies as a function of E, phi and theta in data and simulation with a point-to-point precision of at least 5%.
  2. Measure systematics of pi0/eta mass calibration as a function of detector position to a precision of at least 5 MeV.
  3. Demonstrate agreement of photon reconstruction efficiency and resolution between data and simulation as a function of E, phi and theta to within 5%.

Action Items

  1. Short term CALWG Items
  2. Long term items
  3. Calorimeter work packages
  4. AI - BCAL Monitoring:
    1. Hydra Real-Time Dashboard
    2. Hydra 24-hr Monitoring Logs

Tentative Agenda

  1. Announcements
  2. Past Meeting Minutes
  3. Shutdown tasks
    1. Review Action Items
  4. Run Operations
    1. FCAL Photon energy range --ig
    2. BCAL
    3. CCAL
  5. Calibration
  6. Monitoring
    1. BCAL (offline)
    2. FCAL (offline)
      • monitoring ver06: Batch 1 done, nothing to report
      • to do list FCAL LED efficiency and timing status:
        • 'TotalInte' (Sum over ADC Integrals showing peaks for LED configurations)
          • Runs with only one, large value, related to entering the dark room to replace bases
          • Noisy run in SRC data? Peaks in-between?
          • Rerun LED timing status after calibration is final (started for PrimeX)
        • LED hi rate issue
    3. CCAL (offline)
      • monitoring ver06: Batch 1 done, nothing to report
  7. Reconstruction
  8. Simulations
  9. Any other business


Attendees: Mark, Varun, Elton, Stjepan, Igal, Karthik, Jon, Susan, Malte, Zisis

  1. Announcements: APS is taking place this week. (Tolga and Churamani are there).
  2. Past Meeting Minutes: business arising is dealt with below.
  3. Shutdown tasks
    1. Review Action Items: Karthik reported that Work Package 2 will be wrapped up, by adding conclusions to DocDB-4139.
  4. Run Operations
    1. FCAL
      1. Igal presented a proposal to lower the FCAL range from 2V to 1V which would allow lowering the voltage by ~100V. He did a simulation of the threshold effect.
      2. Since GlueX cares about the lowest energy photons we asked him to show a zoom of the simulation plots to the low energy area.
      3. We also discussed issues with the FCAL energy correction and Igal make recommendations for GlueX energy correction.
      4. We discussed changing the V range just for PrimeX. Malte will look at a board to see if it can be done without uncabling.
    2. FCAL hardware
      1. Mark explained the next steps for FCAL hardware.
        1. Test to see what the effect of uniforming the torque on the the bases to 15 inch ounces would be. Torque whole detector if useful.
        2. Change about 40-50 PMTs to bring up the lowest gain ones.
          1. Test all spare PMTs
          2. Determine which PMTs need to be replaced from pi0 measured gain.
      2. We discussed going to low threshold for CPP. Mark will talk to Matt about exsiting threshold in readout.
    3. BCAL
      1. The BCAL hot channels were discussed. Mark will talk to the AIEC folks about what we can benefit from their work.
    4. CCAL
      1. The CCAL will not be used for CCP. For PrimeX the Muon detector will be removed and the CCAL installed.
  5. Calibration
    1. Stjepan and Karthik working on summary document for PrimeX and SRC.
    2. Stjepan has implemented gain ratio from SRC. Many iterations necessary. More will be needed after next monitoring launch.
  6. Monitoring
    1. Susan presented a TODO list for the FCAL. The CCAL has no problems.